OpenStudy (anonymous):

Solve 3(x + 1) + 6 = 33. 10 8 8.7 -8 I picked 8.7 because when I did the work I got 8.67 am I right?

OpenStudy (anonymous):

multiply the 3(x+1) try to do it i will help u

OpenStudy (anonymous):

Would the 3X + 1 become 4x?

OpenStudy (anonymous):

no it would be 3x+3

OpenStudy (anonymous):

now it will be: 3x+3+6=33

OpenStudy (anonymous):

add 3 and 6

OpenStudy (anonymous):

Okay so I add 3 to every part of the equation? Is that why it's 3x + 3?

OpenStudy (anonymous):

Nows its: 3x + 9 = 33 ?

OpenStudy (anonymous):

33-9 = 24 24 / 3 = 8 is that right?

OpenStudy (anonymous):

yep its 8

OpenStudy (anonymous):

Can I ask you for help with one more? Thank so much for your help already

OpenStudy (anonymous):

yep suuuuure:)

OpenStudy (anonymous):

Solve for x: 1 over 4 (2x − 14) = 4 15 15 over 2 9 11

OpenStudy (anonymous):

first multiply 4(2x-14)

OpenStudy (anonymous):

4 X 2 = 8 and 4 X 14 = 56

OpenStudy (anonymous):

yep so it will be 1 over (8x-56)=4

OpenStudy (anonymous):

now multiply 4 (8x-56)

OpenStudy (anonymous):

4 + 56 = 60 60 / 8 = 7.5

OpenStudy (anonymous):



OpenStudy (anonymous):

4(8x-56)=32x-224

OpenStudy (anonymous):

I have no idea what just happened, out of the choices: 15 15 over 2 9 11 which did you say

OpenStudy (anonymous):

are u sure that these are the chosies?

OpenStudy (anonymous):

Yes I picked 9 and it was wrong.

OpenStudy (anonymous):

can u check the quetion plz?

OpenStudy (anonymous):

coz i got 7 it did it several times and i got 7

OpenStudy (anonymous):

hallo r u still there?